OpenStudy (anonymous):

What is the value of the coefficient "b" when the quadratic equation y = (3x − 5)(2x + 3) is written in standard form? 19 6 −15 −1

OpenStudy (anonymous):

you need to explicitly multiply out the two terms in brackets: (a+b)(c+d) = a*c + a*d + b*c + b*d sooo...: 3x*2x + 3x*3 + (-5)*2x + (-5)*3 = 6x^2 + 9x - 10 x - 15 = 6x^2 - x - 15 .. that is the standard form. ax^2 + bx + c so factor b is (-1).